Problem 1: A corporation will pay a $1.00 dividend (D1) in the next 12 months on a share of common stock. The required rate of return is 5% and the constant growth rate is 4%. Compute the theoretical stock price. Problem 2: A corporation expects to pay dividends (D1) of $1.75 per share at the end of the current year and the current price of its common stock is $30 per share. The expected growth rate is 3.5% and flotation costs of $1.00 per share are anticipated. Making use of the dividend capitalization model, compute the cost of this new common equity.
Answer #1

Problem 1:

Stock Price = Expected Dividend/(Required return - Growth rate)

= 1/(5%-4%)

= $100

Problem 2:

30-1 = 1.75/(Cost of new equity - 3.5%)

Cost of New Equity = 9.53%